### v1

@@ Acknowledgements
Extended Data Fig. 2: The cumulative distribution of the difference between the kinematic ages and tidal evolution timescales $t-t_{\rm tide}$. The results from the observed and synthetic hot Jupiter sample that can be classified as ‘Late-arrived’ are plotted as solid black line and dashed red line, respectively. We print the two-sample KS test $p$-value. The solid orange line denotes the arrival timescale of hot Jupiters via the secular models derived from simulations [31].
Extended Data Fig. 3: The evolutionary patterns of hot Jupiters formed via different origin models. Top panels: The number ratio of the formed hot Jupiters before $t$ over the formed hot Jupiters of all times as a function of age for different origin models. The green and origin lines represent ‘Early ’model and ‘Late’ model, respectively. Bottom panels: The number ratio of the left (formed - tidally disrupted) hot Jupiters before $t$ over the formed hot Jupiters of all times as a function of age for different origin models. The initial conditions are set as the standard case, i.e., the a-distribution of hot Jupiters is set as the results inferred from Kepler data [82] and the initial planetary mass is set as that of cold Jupiters. The modified stellar tidal quality factor $Q^{{}^{\prime}}_{*}=10^{6}$.
@@
Standard case: (a) ‘Early’ model: initial mass and radius distributions as observed warm/cold Jupiters; initial period distribution as the result from Kepler data [82]; (b) ‘Late’ model: initial distributions referring to Hamers et al. 2017 [31].
@@
From theory, the initial mass distribution of hot Jupiters is expected to be similar to that of warm/cold Jupiters if disk migration is dominating. In contrast, under high-eccentricity migration (e.g., secular chaos), hot Jupiters tend to have smaller mass distributions, as smaller planets are more easily driven into high-eccentricity orbits by interactions with other planets [10, 11]. Therefore, the intrinsic initial mass distribution of hot Jupiters is hard to be estimated accurately and is expected to lie between those of observed hot and warm/cold Jupiters.
@@
    the uniform distribution in logarithmic space inferred from early Doppler survey [99];
@@
    the distribution as the disk migration component derived from Nelson et al. (2017) [100], $\frac{{\rm d}N}{{\rm d}x}\propto x^{\gamma-1},\ x\equiv\frac{a}{a_{\rm R}},\ \gamma=-0.04$;
@@
    the distribution as the high-eccentricity migration component derived from Nelson et al. (2017) [100] which have a $\gamma=-1.38$.
@@
    Gyrochronology age. Gyrochronology estimates the age of a main-sequence star based on its rotation period [101, 102]. By adopting the method provided by Spada & Lanzafame [103], we calculate gyrochronology ages for 31,860 Kepler stars based on the rotation data from [104]. Then we select Sun-like stars with $T_{\rm eff}$ in the range of 470065004700-6500 K and remove potential binaries by excluding stars with re-normalized unit weight error (RUWE)>1.2>1.2, leaving 22,570 stars. For the planetary sample, we cross-match with Kepler DR25 catalog and yield 16 hot Jupiters.
    To obtain the age-frequency trend, we divide the stellar sample into 5 bins with equal size according to their gyrochronology age and calculate the frequencies of hot Jupiters by correcting the geometric effect/detection efficiency and eliminating the effect of $\rm[Fe/H]$ with same procedure described in § 3.2 of Chen2023 [21]. Figure S11 shows the frequency of hot Jupiters as a function of gyrochronology age (pink points). As can be seen, $F_{\rm HJ}$ remains relatively stable when $t\lesssim 2$ Gyr and then declines sharply. It is also worth noting that the ages of some stars in the last two bins may be underestimated due to stalled spin-down.
@@
    Isochrone age. [105] provided isotropic ages for 186,301 Kepler stars with a typical uncertainty of $\sim 56\%$ by fitting the isochrone grid. Here we select Sun-like stars with reliable age estimations by applying the following criteria: $\rm GOF>0.99$ & $TAMS<20$ Gyr & $t<14$ Gyr & $\frac{{\rm err}\_t}{t}<1/3$, yielding 17665 Kepler stars. For the planetary sample, we cross-match with the Kepler DR25 catalog and obtained 10 hot Jupiters. We then divided the stellar sample into 5 bins and derive the frequency of hot Jupiters of different age bins. As shown in Figure S11 (pink points), $F_{\rm HJ}$ generally decreases when $t>2$ Gyr.
@@
In the above analyses, we fit the observed frequency-age trend of hot Jupiters to theoretical predictions, assuming a common tidal quality factor. Some previous studies have suggested that $Q^{{}^{\prime}}_{*}$ depend on factors such as stellar mass, age, evolutionary stage, and the orbital periods of hot Jupiters [32, 24, 33]. Thus, in this subsection, we explore the influence of the above dependences of $Q^{{}^{\prime}}_{*}$ as follows:
@@
    Stellar properties. $Q^{{}^{\prime}}$ may decrease at specific stellar evolutionary stages (e.g., subgiant) due to the change of stellar structures [e.g., the increase of stellar radii and convective envelopes; 106, 32]. In this work, the stellar sample has been restricted to Sun-like stars, i.e., main-sequence stars with $T_{\rm eff}$ in the range of 4700-6500 K. In Figure S12, we compare the mass distributions for host stars of different kinematic ages, which are statistically indistinguishable (with KS $p$-values >0.1>0.1). Moreover, $Q^{{}^{\prime}}_{*}$ does not differ significantly with age during the main-sequence stage [see fig. 3 of 33]. Therefore, in this work, it is reasonable to fit using a common tidal factor when considering only stellar properties.
@@
    Planetary orbital period (tidal forcing period). As suggested by the observational [24] and theoretical analyses [33], the stellar $Q^{{}^{\prime}}_{*}$ may vary with the orbital period of hot Jupiters $P$, with an empirical formula,
@@
    where $P_{\rm tide}$ and $P_{\rm spin}$ are the tidal force period and stellar spin period, respectively. In the top panel of Figure S13, we show the derived $Q^{*}_{*}$ as a function of the orbital period $P$, where $P_{\rm spin}$ is set as a typical value of 18 days for Sun-like stars [101, 103]. As can be seen, $Q^{{}^{\prime}}_{*}$ remains constant when $P\gtrsim 4$ days, while for $P\lesssim 3$ day, $Q^{{}^{\prime}}_{*}$ increases with decreasing $P$. We also calculate the in-sprial timescale $t_{\rm in}$. The stellar and planetary periods are taken as the median values for the ‘Early’ model population in the standard case. As shown in the bottom panel, $t_{\rm in}$ become longer but does not differ significantly compared to typical age uncertainties (i.e., 0.2-2 Gyrs when t $\sim 0.5-10$ Gyr). This is not unexpected since $t_{\rm in}\propto P^{6.5}$ , while the variation in $Q^{{}^{\prime}}_{*}$ primarily occurs at shorter periods (corresponding to $t_{\rm in}\lesssim 1$ Gyr) with a weaker correlation, scaling as $P^{-3.1}$.
@@
    It is worth noting that the best fit of the hybrid model obtained using $Q^{{}^{\prime}}_{*}$ as [24] does not align very well with observations, as the inflection point appears earlier than 1 Gyr. Moreover, the best fit of Hybrid model obtained using $Q^{{}^{\prime}}_{*}$ as [24] exhibits a poorer agreement with observations compared to that obtained using a common $Q^{{}^{\prime}}_{*}$ (Extended Data Fig. 6), with a larger AIC scores of -105.6 ($\rm\Delta AIC=10.9$).
@@
The formation timescale of hot Jupiter via the secular chaos mechanism depends on the initial conditions of the planetary system [10], which remain unclear. To test how sensitive their conclusion is to this choice of secular chaos model and its initial conditions, we also adopted the secular chaos model from Teyssandier et al. (2019) [34] as the ‘late’ model and fitted it to the observational age-frequency trend.
@@
We then perform numerical simulations for the tidal decays of hot Jupiters with various stellar tidal quality factor ($Q^{{}^{\prime}}_{*}$). Figure S16 shows the best match of $F_{\rm HJ}$ from the hybrid model comparing to the observational results. The Hybrid model is preferred comparing to the single ‘Early’ model and single ‘Late’ model with confidence levels of $\gtrsim 2$-$\sigma$, respectively. The best-fits for the $f_{\rm Late}$ and $Q^{{}^{\prime}}_{*}$ are $45^{+18}_{-16}\%$ and $5.6^{+0.2}_{-0.3}$, which are nearly identical to those when adopting the secular model from Hamers et al. (2017) [31] though with a larger AIC score of -111.9.
@@
Figure S13: Top: The stellar tidal factor $Q^{{}^{\prime}}_{*}$ (Top) as a function of orbital period ($P$) referring to [24]. Bottom: The in-sprial timescales $t_{\rm in}$ as a function of $P$ by considering the above $Q^{{}^{\prime}}_{*}-P$ dependence. The stellar and planetary periods are taken as the median values for the ‘Early’ model population in the standard case. For comparison, we also plot the result when taking a constant $\log Q^{{}^{\prime}}_{*}=5.0$.
Figure S14: The number ratio of the left (formed - tidally disrupted) hot Jupiters before $t$ over the formed hot Jupiters of all times as a function of age for the ‘Early’ model (Top) and ‘Late’ model (Bottom), respectively. The initial conditions are set as the standard case, i.e., the a-distribution of hot Jupiters is set as the results inferred from Kepler data [82] and the initial planetary mass is set as that of cold Jupiters. The modified stellar tidal quality factor $Q^{{}^{\prime}}_{*}$ is set as Equation S1. For comparison, we also plot the result when taking a constant $\log Q^{{}^{\prime}}_{*}=5.0$.
Figure S15: Similar to Extended Data Fig. 6 but the $Q^{{}^{\prime}}_{*}$ is set as the results derived in [24], which varies with orbital period (Eq. S1). The right panel displays the relative likelihood in logarithm as a function of $f_{\rm Late}$.
Figure S16: Similar to Extended Data Fig. 6 but the ‘late’ model is adopt as the secular chaos model from [34].

### v2

@@ Acknowledgements
Extended Data Fig. 2: The cumulative distribution of the difference between the kinematic ages and tidal evolution timescales $t-t_{\rm tide}$. The results from the observed and synthetic hot Jupiter sample that can be classified as ‘Late-arrived’ are plotted as solid black line and dashed red line, respectively. We print the two-sample KS test $p$-value. The solid orange line denotes the arrival timescale of hot Jupiters via the secular models derived from simulations [2017MNRAS.464..688H].
Extended Data Fig. 3: The evolutionary patterns of hot Jupiters formed via different origin models. Top panels: The number ratio of the formed hot Jupiters before $t$ over the formed hot Jupiters of all times as a function of age for different origin models. The green and origin lines represent ‘Early ’model and ‘Late’ model, respectively. Bottom panels: The number ratio of the left (formed - tidally disrupted) hot Jupiters before $t$ over the formed hot Jupiters of all times as a function of age for different origin models. The initial conditions are set as the standard case, i.e., the a-distribution of hot Jupiters is set as the results inferred from Kepler data [2016A&A...587A..64S] and the initial planetary mass is set as that of cold Jupiters. The modified stellar tidal quality factor $Q^{{}^{\prime}}_{*}=10^{6}$.
@@
Standard case: (a) ‘Early’ model: initial mass and radius distributions as observed warm/cold Jupiters; initial period distribution as the result from Kepler data [2016A&A...587A..64S]; (b) ‘Late’ model: initial distributions referring to Hamers et al. 2017 [2017MNRAS.464..688H].
@@
From theory, the initial mass distribution of hot Jupiters is expected to be similar to that of warm/cold Jupiters if disk migration is dominating. In contrast, under high-eccentricity migration (e.g., secular chaos), hot Jupiters tend to have smaller mass distributions, as smaller planets are more easily driven into high-eccentricity orbits by interactions with other planets [2011ApJ...735..109W, 2015ApJ...805...75P]. Therefore, the intrinsic initial mass distribution of hot Jupiters is hard to be estimated accurately and is expected to lie between those of observed hot and warm/cold Jupiters.
@@
    the uniform distribution in logarithmic space inferred from early Doppler survey [2007ARA&A..45..397U];
@@
    the distribution as the disk migration component derived from Nelson et al. (2017) [2017AJ....154..106N], $\frac{{\rm d}N}{{\rm d}x}\propto x^{\gamma-1},\ x\equiv\frac{a}{a_{\rm R}},\ \gamma=-0.04$;
@@
    the distribution as the high-eccentricity migration component derived from Nelson et al. (2017) [2017AJ....154..106N] which have a $\gamma=-1.38$.
@@
    Gyrochronology age. Gyrochronology estimates the age of a main-sequence star based on its rotation period [2010ApJ...722.. [^2] B, 2010ApJ...721..675B]. By adopting the method provided by Spada & Lanzafame [2014ApJS..211...24M], we calculate gyrochronology ages for 31,860 Kepler stars based on the rotation data from [2020A&A...636A..76S]. Then we select Sun-like stars with $T_{\rm eff}$ in the range of 470065004700-6500 K and remove potential binaries by excluding stars with re-normalized unit weight error (RUWE)>1.2>1.2, leaving 22,570 stars. For the planetary sample, we cross-match with Kepler DR25 catalog and yield 16 hot Jupiters.
    To obtain the age-frequency trend, we divide the stellar sample into 5 bins with equal size according to their gyrochronology age and calculate the frequencies of hot Jupiters by correcting the geometric effect/detection efficiency and eliminating the effect of $\rm[Fe/H]$ with same procedure described in § 3.2 of Chen2023 [2023PNAS..12004179C]. Figure S11 shows the frequency of hot Jupiters as a function of gyrochronology age (pink points). As can be seen, $F_{\rm HJ}$ remains relatively stable when $t\lesssim 2$ Gyr and then declines sharply. It is also worth noting that the ages of some stars in the last two bins may be underestimated due to stalled spin-down.
@@
    Isochrone age. [2020AJ....159..280B] provided isotropic ages for 186,301 Kepler stars with a typical uncertainty of $\sim 56\%$ by fitting the isochrone grid. Here we select Sun-like stars with reliable age estimations by applying the following criteria: $\rm GOF>0.99$ & $TAMS<20$ Gyr & $t<14$ Gyr & $\frac{{\rm err}\_t}{t}<1/3$, yielding 17665 Kepler stars. For the planetary sample, we cross-match with the Kepler DR25 catalog and obtained 10 hot Jupiters. We then divided the stellar sample into 5 bins and derive the frequency of hot Jupiters of different age bins. As shown in Figure S11 (pink points), $F_{\rm HJ}$ generally decreases when $t>2$ Gyr.
@@
In the above analyses, we fit the observed frequency-age trend of hot Jupiters to theoretical predictions, assuming a common tidal quality factor. Some previous studies have suggested that $Q^{{}^{\prime}}_{*}$ depend on factors such as stellar mass, age, evolutionary stage, and the orbital periods of hot Jupiters [2017ApJ...849L..11W, 2018AJ....155..165P, 2020MNRAS.498.2270B]. Thus, in this subsection, we explore the influence of the above dependences of $Q^{{}^{\prime}}_{*}$ as follows:
@@
    Stellar properties. $Q^{{}^{\prime}}$ may decrease at specific stellar evolutionary stages (e.g., subgiant) due to the change of stellar structures [e.g., the increase of stellar radii and convective envelopes; 2009ApJ...705L..81V, 2017ApJ...849L..11W]. In this work, the stellar sample has been restricted to Sun-like stars, i.e., main-sequence stars with $T_{\rm eff}$ in the range of 4700-6500 K. In Figure S12, we compare the mass distributions for host stars of different kinematic ages, which are statistically indistinguishable (with KS $p$-values >0.1>0.1). Moreover, $Q^{{}^{\prime}}_{*}$ does not differ significantly with age during the main-sequence stage [see fig. 3 of 2020MNRAS.498.2270B]. Therefore, in this work, it is reasonable to fit using a common tidal factor when considering only stellar properties.
@@
    Planetary orbital period (tidal forcing period). As suggested by the observational [2018AJ....155..165P] and theoretical analyses [2020MNRAS.498.2270B], the stellar $Q^{{}^{\prime}}_{*}$ may vary with the orbital period of hot Jupiters $P$, with an empirical formula,
@@
    where $P_{\rm tide}$ and $P_{\rm spin}$ are the tidal force period and stellar spin period, respectively. In the top panel of Figure S13, we show the derived $Q^{*}_{*}$ as a function of the orbital period $P$, where $P_{\rm spin}$ is set as a typical value of 18 days for Sun-like stars [2010ApJ...722.. [^2] B, 2014ApJS..211...24M]. As can be seen, $Q^{{}^{\prime}}_{*}$ remains constant when $P\gtrsim 4$ days, while for $P\lesssim 3$ day, $Q^{{}^{\prime}}_{*}$ increases with decreasing $P$. We also calculate the in-sprial timescale $t_{\rm in}$. The stellar and planetary periods are taken as the median values for the ‘Early’ model population in the standard case. As shown in the bottom panel, $t_{\rm in}$ become longer but does not differ significantly compared to typical age uncertainties (i.e., 0.2-2 Gyrs when t $\sim 0.5-10$ Gyr). This is not unexpected since $t_{\rm in}\propto P^{6.5}$ , while the variation in $Q^{{}^{\prime}}_{*}$ primarily occurs at shorter periods (corresponding to $t_{\rm in}\lesssim 1$ Gyr) with a weaker correlation, scaling as $P^{-3.1}$.
@@
    It is worth noting that the best fit of the hybrid model obtained using $Q^{{}^{\prime}}_{*}$ as [2018AJ....155..165P] does not align very well with observations, as the inflection point appears earlier than 1 Gyr. Moreover, the best fit of Hybrid model obtained using $Q^{{}^{\prime}}_{*}$ as [2018AJ....155..165P] exhibits a poorer agreement with observations compared to that obtained using a common $Q^{{}^{\prime}}_{*}$ (Extended Data Fig. 6), with a larger AIC scores of -105.6 ($\rm\Delta AIC=10.9$).
@@
The formation timescale of hot Jupiter via the secular chaos mechanism depends on the initial conditions of the planetary system [2011ApJ...735..109W], which remain unclear. To test how sensitive their conclusion is to this choice of secular chaos model and its initial conditions, we also adopted the secular chaos model from Teyssandier et al. (2019) [2019MNRAS.486.2265T] as the ‘late’ model and fitted it to the observational age-frequency trend.
@@
We then perform numerical simulations for the tidal decays of hot Jupiters with various stellar tidal quality factor ($Q^{{}^{\prime}}_{*}$). Figure S16 shows the best match of $F_{\rm HJ}$ from the hybrid model comparing to the observational results. The Hybrid model is preferred comparing to the single ‘Early’ model and single ‘Late’ model with confidence levels of $\gtrsim 2$-$\sigma$, respectively. The best-fits for the $f_{\rm Late}$ and $Q^{{}^{\prime}}_{*}$ are $45^{+18}_{-16}\%$ and $5.6^{+0.2}_{-0.3}$, which are nearly identical to those when adopting the secular model from Hamers et al. (2017) [2017MNRAS.464..688H] though with a larger AIC score of -111.9.
@@
Figure S13: Top: The stellar tidal factor $Q^{{}^{\prime}}_{*}$ (Top) as a function of orbital period ($P$) referring to [2018AJ....155..165P]. Bottom: The in-sprial timescales $t_{\rm in}$ as a function of $P$ by considering the above $Q^{{}^{\prime}}_{*}-P$ dependence. The stellar and planetary periods are taken as the median values for the ‘Early’ model population in the standard case. For comparison, we also plot the result when taking a constant $\log Q^{{}^{\prime}}_{*}=5.0$.
Figure S14: The number ratio of the left (formed - tidally disrupted) hot Jupiters before $t$ over the formed hot Jupiters of all times as a function of age for the ‘Early’ model (Top) and ‘Late’ model (Bottom), respectively. The initial conditions are set as the standard case, i.e., the a-distribution of hot Jupiters is set as the results inferred from Kepler data [2016A&A...587A..64S] and the initial planetary mass is set as that of cold Jupiters. The modified stellar tidal quality factor $Q^{{}^{\prime}}_{*}$ is set as Equation S1. For comparison, we also plot the result when taking a constant $\log Q^{{}^{\prime}}_{*}=5.0$.
Figure S15: Similar to Extended Data Fig. 6 but the $Q^{{}^{\prime}}_{*}$ is set as the results derived in [2018AJ....155..165P], which varies with orbital period (Eq. S1). The right panel displays the relative likelihood in logarithm as a function of $f_{\rm Late}$.
Figure S16: Similar to Extended Data Fig. 6 but the ‘late’ model is adopt as the secular chaos model from [2019MNRAS.486.2265T].